OpenStudy (howard-wolowitz):

Can someone help me with this question:

OpenStudy (calculusxy):

Most likely the quotient property, I believe.

OpenStudy (howard-wolowitz):

hmm ok I dont know at all

OpenStudy (calculusxy):

"The logarithm of a product property says log2 8a = log2 8 + log2 a, and log2 8 = 3. You can use the similarity between the properties of exponents and logarithms to find the property for the logarithm of a quotient. With exponents, to multiply two numbers with the same base, you add the exponents," (montereyinstitute.org).

OpenStudy (calculusxy):

The \[\log_{4} \] remains the same, but the 7 and the 2 are multiplied to get 14. \[\log_{4} 14\]

Nnesha (nnesha):

yep right log properties! product rule \[\huge\rm log_b x + \log_b y = \log_b( x \times y )\] addition ----> multiplication

OpenStudy (calculusxy):

Exactly! @Nnesha

OpenStudy (calculusxy):

@Howard-Wolowitz Got it?

OpenStudy (howard-wolowitz):

yes I do

OpenStudy (howard-wolowitz):

actually so its A and not B

OpenStudy (howard-wolowitz):

because it the product property

Nnesha (nnesha):

hmm you don't have to solve it ?? :O

Nnesha (nnesha):

my bad o^_^o

OpenStudy (calculusxy):

Should have understood! UGH! Really sorry

OpenStudy (howard-wolowitz):

me sorry, me friends, I got it so its actually Quotient property.......many thanks amigo

OpenStudy (howard-wolowitz):

BEcause its that ^

Nnesha (nnesha):

how is it quotient property ?

OpenStudy (howard-wolowitz):

me confused. I thought you said you thought it was C or A... but I thought it was D

Nnesha (nnesha):

just look at the log properties

OpenStudy (howard-wolowitz):

yes product

Nnesha (nnesha):

quotient rule\[\huge\rm log_b y - \log_b x = \log_b \frac{ x }{ y}\] to condense you can change subtraction to division product rule \[\huge\rm log_b x + \log_b y = \log_b( x \times y )\] addition ----> multiplication power rule \[\huge\rm log_b x^y = y \log_b x\] CHANGE OF BASE FORMULA: change of base formula \[\huge\rm log_a b = \frac{ \log b }{ \log a }\] here are log properties write it down on ur notes! you need these

OpenStudy (howard-wolowitz):

ok many thanks BIG GURL .. me amigo

OpenStudy (howard-wolowitz):

ill go with product

Nnesha (nnesha):

bec there is a plus sign you would convert that equation to multiplication so yes you should go with product!rule

OpenStudy (howard-wolowitz):

MANY , many thanks me BIG GURLS ... I happy

OpenStudy (anonymous):

He is a 45 year old civics teacher online and his real name is Gregory Notowitz

Nnesha (nnesha):

np hope you'll familiar with these when u make a new post!

OpenStudy (howard-wolowitz):

ok BIG GURL